Math 121 College Algebra Test F 7. A ball is thrown directly upward from a height of 6 ft with an initial velocity of 20 fu/sec. a) Express the height s of the ball, in feet, seconds after it has been thrown. b) Determine the time at which the ball reaches its maximum height. Round the answer to two decimals c) Find the maximum height of the ball. Round the answer to two decimals
Answer #1

7. (a) yi = 6 ft

v0 =20 ft/s

a = - 32.17 ft/s^2


yf - yi = v0 t + a t^2 / 2

h - 6 = (20 t) + (-32.17 t^2 / 2)

h = 6 + 20t - 16 t^2

(B) h will be max when dh/dt = 0


20 - 32 t = 0

t = 0.62 sec

(C) putting value of t,

h = 12.22 ft